# Divergence Minimization Preference Optimization for Diffusion Model Alignment

### v1

@@ B.2 Details of Datasets
Pick-a-Pic [14]: This comprehensive dataset captures real user preferences from the Pick-a-Pic web platform, where users generate images from text prompts. The collection encompasses more than 500,000 preference examples derived from over 35,000 unique prompts. Each entry consists of a text prompt, two AI-generated images, and user feedback indicating their preferred image or marking cases where no clear preference exists. The dataset incorporates outputs from various generative models, including Stable Diffusion 2.1, Dreamlike Photoreal 2.05, and multiple Stable Diffusion XL configurations, utilizing different classifier-free guidance parameters during generation.
Parti-Prompts [41]: This benchmark dataset features more than 1,600 carefully crafted English prompts designed to evaluate text-to-image model capabilities. The prompts cover diverse categories and present varied challenges, enabling comprehensive assessment of model performance across multiple evaluation criteria.
HPS V2 [37]: This preference-based collection contains 98,807 generated images sourced from 25,205 distinct prompts. The dataset structure allows for multiple image generations per prompt, with users selecting their preferred output while other variations serve as negative examples. The distribution includes 23,722 prompts with four associated images, 953 prompts with three images, and 530 prompts paired with two images.
TEd-bench [12]: Textual Editing Benchmark is a text-to-image editing benchmark that provides pairs of real images and corresponding editing prompts. For example, given a source image of a dog running, the target prompt might be “A cat running on the grass”, in which case the model is expected to modify the original image so that it depicts the desired semantics. We evaluate both existing baselines and our proposed method on TEd-bench by applying an automated reward model to score each edited image, and report the average reward as our editing performance metric.
InstructPix2Pix [3]: This specialized collection focuses on instruction-based image editing capabilities. The dataset enables models to modify existing images according to natural language instructions, such as "make the clouds rainy." The system processes both the editing instruction and the original image to produce the desired modifications. Our experimental evaluation utilized 1,000 test samples from this dataset to assess instruction-following image editing performance. Note that although reward models such as CLIP [26] may not comprehensively capture the semantics of instruction-style prompts when scoring edited images, we maintain that if the key terms from the instruction appear faithfully in the generated image, this constitutes a valid comparison.
@@ C.1 More Qualitative Results
To further demonstrate the effectiveness of DMPO, we provide additional qualitative results on text-to-image alignment and image editing tasks. These results in Figure 5, Figure 6, Figure 7, and Figure 8 clearly show that DMPO consistently produces superior outputs compared to baseline methods. We list the prompts used in the Figures as follows:
Prompts used in Figure 5:
  • a painting of a fox in the style of starry night.
  • Cyberpunk cat.
  • The image is a vibrant and intricate illustration of a man, with a focus on his shoulder and head, created using inkpen and Unreal Engine technology.
  • A cat with two horns on its head.
  • A spoon dressed up with eyes and a smile.
  • A papaya fruit dressed as a sailor.
  • A giant dinosaur frozen into a glacier and recently discovered by scientists, cinematic still.
  • photo of a zebra dressed suit and tie sitting at a table in a bar with a bar stools, award winning photography.
Prompts used in Figure 6:
  • an image of a photo model wearing a red lace dress, standing in the jungle.
  • a child in the air while jumping on a trampoline.
  • Two cups of coffee, one with latte art of the words "LOVE" written in one. The other has latte art of the words "PEACE" written in the other.
  • anthropomorphic coffee bean drinking coffee.
  • 16-year-old teenager wearing a white bear-ear hat with a smirk on their face.
  • Frontal portrait of an anime girl with pink hair and sunglasses wearing a white t-shirt.
  • A giant dinosaur frozen into a glacier and recently discovered by scientists, cinematic still.
  • Of the lunar module landing on a hydrogen lake on Titan, through a foggy yellow smog.
Prompts used in Figure 7:
  • A girl with pink pigtails and face tattoos.
  • A cat with two horns on its head.
  • A hand-drawn cute gnome holding a pumpkin in an autumn disguise, portrayed in a detailed close-up of the face with warm lighting and high detail.
  • A cute puppy leading a session of the United Nations, newspaper photography.
  • A purple raven flying over Big Sur, light fog, deep focus+closeup, hyper-realistic, volumetric lighting, dramatic lighting, beautiful composition, intricate details, instagram, trending, photograph, film grain and noise, 8K, cinematic, post-production.
  • photo of a zebra dressed suit and tie sitting at a table in a bar with a bar stools, award winning photography.
Prompts used in Figure 8:
  • A photo of a bird spreading wings.
  • A half eaten pizza.
  • A jumping horse.
  • A goat and a cat hugging.
  • A teddy bear holding a cup.
  • Image of a cat wearing a floral shirt.
  • A door with a pet entrance.
  • A photo of a cat wearing a hat.
@@ C.2 More Quantitative Results
We also conduct quantitative evaluations and analyses for all experiments based on SD1.5 and SDXL. As shown in Table 3, Table 4, Table 5 and Table 6, DMPO demonstrates strong generalization ability, consistently outperforming other baselines in both reward model scores and win rates across both base models.
Figure 5: Images generated by different models (based on SD1.5) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2. Prompts used can be found in Section C.1
Figure 6: Images generated by different models (based on SD1.5) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2. Prompts used can be found in Section C.1.
Figure 7: Images generated by different models (based on SDXL) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2. Prompts used can be found in Section C.1.
Figure 8: Images edited by different models (based on SD1.5) for various prompts which are selected from TEd-bench. Prompts used can be found in Section C.1.
@@
Pick-a-Pic V2 SDXL 0.2203 0.2661 0.3609 5.9892 0.5111
DMPO 0.2264 0.2730 0.3741 5.9422 0.8563
SFT 0.2224 0.2675 0.3624 5.9239 0.5834
Diff-DPO 0.2256 0.2709 0.3722 5.9890 0.7584
MAPO 0.2213 0.2682 0.3615 6.1196 0.6226
DSPO 0.2256 0.2684 0.3615 5.9598 0.6831
HPS V2 SDXL 0.2271 0.2730 0.3775 6.1125 0.6749
DMPO 0.2320 0.2804 0.3914 6.1883 1.0169
SFT 0.2277 0.2762 0.3784 6.0638 0.6998
Diff-DPO 0.2314 0.2777 0.3890 6.1546 0.9610
MAPO 0.2279 0.2765 0.3801 6.2134 0.7839
DSPO 0.2285 0.2754 0.3795 6.0545 0.7385
Parti Prompt SDXL 0.2249 0.2714 0.3551 5.7648 0.6010
DMPO 0.2290 0.2780 0.3769 5.8598 1.0399
SFT 0.2257 0.2701 0.3531 5.7239 0.6953
Diff-DPO 0.2286 0.2763 0.3688 5.7900 0.9515
MAPO 0.2250 0.2732 0.3561 5.9089 0.7031
DSPO 0.2256 0.2714 0.3596 5.7598 0.7600
Pick-a-Pic V2 DMPO SDXL 75.20 80.20 61.60 48.00 69.80
SFT SDXL 54.20 67.50 53.50 43.20 61.80
Diff-DPO SDXL 74.50 79.00 61.80 51.00 69.00
MAPO SDXL 55.60 68.00 51.20 66.60 64.20
DSPO SDXL 56.30 69.20 52.60 44.80 60.50
DMPO SFT 65.80 72.00 60.20 53.20 66.20
DMPO Diff-DPO 51.50 67.00 50.80 45.40 61.20
DMPO MAPO 66.20 74.00 61.60 35.80 64.60
DMPO DSPO 63.40 73.40 63.40 50.30 63.20
HPS V2 DMPO SDXL 67.75 87.25 63.75 47.75 71.25
SFT SDXL 54.75 70.00 53.00 61.25 60.00
Diff-DPO SDXL 70.50 81.25 61.75 59.25 68.75
MAPO SDXL 54.00 79.50 52.00 70.00 62.75
DSPO SDXL 53.25 71.25 55.25 57.35 64.75
DMPO SFT 60.75 65.50 58.75 50.00 63.50
DMPO Diff-DPO 48.00 72.00 53.00 43.00 52.50
DMPO MAPO 62.25 75.25 59.50 38.50 66.75
DMPO DSPO 63.75 73.00 59.75 48.25 62.25
Parti Prompt DMPO SDXL 68.72 79.96 61.46 51.69 71.75
SFT SDXL 50.44 66.58 50.75 52.14 56.69
Diff-DPO SDXL 66.89 76.89 60.64 54.96 69.30
MAPO SDXL 49.94 67.34 50.80 69.24 59.44
DSPO SDXL 52.44 65.17 53.31 53.22 58.71
DMPO SFT 62.22 66.31 59.23 50.69 53.46
DMPO Diff-DPO 51.07 61.89 52.77 45.62 52.14
DMPO MAPO 64.22 70.22 62.01 44.21 66.24
DMPO DSPO 60.12 63.08 58.25 51.49 57.52

### v2

@@ B.2 Details of Datasets
Pick-a-Pic [Kirstain et al., 2023]: This comprehensive dataset captures real user preferences from the Pick-a-Pic web platform, where users generate images from text prompts. The collection encompasses more than 500,000 preference examples derived from over 35,000 unique prompts. Each entry consists of a text prompt, two AI-generated images, and user feedback indicating their preferred image or marking cases where no clear preference exists. The dataset incorporates outputs from various generative models, including Stable Diffusion 2.1, Dreamlike Photoreal 2.05, and multiple Stable Diffusion XL configurations, utilizing different classifier-free guidance parameters during generation.
Parti-Prompts [Yu et al., 2022]: This benchmark dataset features more than 1,600 carefully crafted English prompts designed to evaluate text-to-image model capabilities. The prompts cover diverse categories and present varied challenges, enabling comprehensive assessment of model performance across multiple evaluation criteria.
HPS V2 [Wu et al., 2023]: This preference-based collection contains 98,807 generated images sourced from 25,205 distinct prompts. The dataset structure allows for multiple image generations per prompt, with users selecting their preferred output while other variations serve as negative examples. The distribution includes 23,722 prompts with four associated images, 953 prompts with three images, and 530 prompts paired with two images.
TEd-bench [Kawar et al., 2023]: Textual Editing Benchmark is a text‐to‐image editing benchmark that provides pairs of real images and corresponding editing prompts. For example, given a source image of a dog running, the target prompt might be “A cat running on the grass”, in which case the model is expected to modify the original image so that it depicts the desired semantics. We evaluate both existing baselines and our proposed method on TEd-bench by applying an automated reward model to score each edited image, and report the average reward as our editing performance metric.
InstructPix2Pix [Brooks et al., 2023]: This specialized collection focuses on instruction-based image editing capabilities. The dataset enables models to modify existing images according to natural language instructions, such as "make the clouds rainy." The system processes both the editing instruction and the original image to produce the desired modifications. Our experimental evaluation utilized 1,000 test samples from this dataset to assess instruction-following image editing performance. Note that although reward models such as CLIP [Radford et al., 2021] may not comprehensively capture the semantics of instruction-style prompts when scoring edited images, we maintain that if the key terms from the instruction appear faithfully in the generated image, this constitutes a valid comparison.
@@ C.1 More Qualitative Results
To further demonstrate the effectiveness of DMPO, we provide additional qualitative results on text-to-image alignment and image editing tasks. These results in Figure˜5, Figure˜6, Figure˜7, and Figure˜8 clearly show that DMPO consistently produces superior outputs compared to baseline methods.
@@ C.2 More Quantitative Results
We also conduct quantitative evaluations and analyses for all experiments based on SD1.5 and SDXL. As shown in Table˜7, Table˜8, Table˜2 and Table˜5, DMPO demonstrates strong generalization ability, consistently outperforming other baselines in both reward model scores and win rates across both base models.
Figure 5: Images generated by different models (based on SD1.5) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2.
Figure 6: Images generated by different models (based on SD1.5) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2.
Figure 7: Images generated by different models (based on SDXL) for various prompts which are selected from Pick-a-Pic V2, Parti-Prompt and HPS V2.
Figure 8: Images edited by different models (based on SD1.5) for various prompts which are selected from TEd-bench.
Pick-a-Pic V2 DMPO SD1.5 82.40 84.80 65.60 75.60 80.40
SFT SD1.5 74.60 82.60 61.60 72.40 77.80
Diff-DPO SD1.5 73.20 70.00 61.20 69.40 64.20
Diff-KTO SD1.5 73.60 84.60 58.60 74.00 80.00
DSPO SD1.5 74.60 84.80 61.60 73.60 76.20
DMPO SFT 67.20 50.60 52.80 53.20 50.20
DMPO Diff-DPO 73.40 77.20 56.80 64.80 68.40
DMPO Diff-KTO 69.00 50.80 55.20 55.20 50.20
DMPO DSPO 68.40 50.80 54.80 52.20 50.80
HPS V2 DMPO SD1.5 85.50 86.50 61.00 78.75 79.75
SFT SD1.5 77.00 89.00 56.00 73.75 81.50
Diff-DPO SD1.5 77.50 69.25 57.75 69.5 64.50
Diff-KTO SD1.5 75.25 89.50 55.00 75.75 79.25
DSPO SD1.5 74.50 89.00 58.00 78.25 78.75
DMPO SFT 67.25 47.50 56.00 57.75 53.50
DMPO Diff-DPO 74.00 74.25 58.00 70.00 72.00
DMPO Diff-KTO 68.00 47.50 56.50 55.50 53.50
DMPO DSPO 65.75 50.00 59.00 57.00 54.00
Parti Prompt DMPO SD1.5 76.72 82.67 62.01 71.69 75.25
SFT SD1.5 66.36 83.88 54.35 71.51 72.37
Diff-DPO SD1.5 67.16 64.86 54.84 60.66 63.11
Diff-KTO SD1.5 65.56 84.38 54.41 69.97 73.86
DSPO SD1.5 67.30 85.17 53.31 70.22 73.71
DMPO SFT 64.64 51.31 57.23 51.69 56.00
DMPO Diff-DPO 69.30 72.54 58.15 65.62 69.55
DMPO Diff-KTO 67.03 48.40 59.13 55.15 55.46
DMPO DSPO 66.12 47.14 59.25 53.49 55.88
Pick-a-Pic V2 DMPO SDXL 75.20 80.20 61.60 48.00 69.80
SFT SDXL 54.20 67.50 53.50 43.20 61.80
Diff-DPO SDXL 74.50 79.00 61.80 51.00 69.00
MAPO SDXL 55.60 68.00 51.20 66.60 64.20
DSPO SDXL 56.30 69.20 52.60 44.80 60.50
DMPO SFT 65.80 72.00 60.20 53.20 66.20
DMPO Diff-DPO 51.50 67.00 50.80 45.40 61.20
DMPO MAPO 66.20 74.00 61.60 35.80 64.60
DMPO DSPO 63.40 73.40 63.40 50.30 63.20
HPS V2 DMPO SDXL 70.75 87.25 63.75 47.75 71.25
SFT SDXL 54.75 70.00 53.00 61.25 60.00
Diff-DPO SDXL 70.50 81.25 61.75 59.25 68.75
MAPO SDXL 54.00 79.50 52.00 70.00 62.75
DSPO SDXL 53.25 71.25 55.25 57.35 64.75
DMPO SFT 60.75 65.50 58.75 50.00 63.50
DMPO Diff-DPO 51.25 72.00 53.00 43.00 52.50
DMPO MAPO 62.25 75.25 59.50 38.50 66.75
DMPO DSPO 63.75 73.00 59.75 48.25 62.25
Parti Prompt DMPO SDXL 68.72 79.96 61.46 51.69 71.75
SFT SDXL 50.44 66.58 50.75 52.14 56.69
Diff-DPO SDXL 66.89 76.89 60.64 54.96 69.30
MAPO SDXL 49.94 67.34 50.80 69.24 59.44
DSPO SDXL 52.44 65.17 53.31 53.22 58.71
DMPO SFT 62.22 66.31 59.23 50.69 53.46
DMPO Diff-DPO 56.07 61.89 52.77 45.62 52.14
DMPO MAPO 64.22 70.22 62.01 44.21 66.24
DMPO DSPO 60.12 63.08 58.25 51.49 57.52
Table 6: (a) Win rate (%) comparisons on Pick-a-Pic V2, HPS V2 and Parti-Prompt datasets for all baselines (fine-tuned on SD3.5-M) versus SD3.5-M. (b) Win rate (%) comparison between DMPO versus other baselines, win rates that surpass 50 % are in green, below 50 % are in red. For simplicity, "Diff" represents "Diffusion".
Pick-a-Pic V2 DMPO SD3.5-M 65.20 70.10 58.40 54.40 59.20
Diff-DPO SD3.5-M 61.50 66.00 53.80 52.00 56.00
DMPO Diff-DPO 54.50 59.00 59.80 53.40 57.20
HPS V2 DMPO SD3.5-M 63.75 77.50 63.75 52.75 68.25
Diff-DPO SD3.5-M 58.50 71.25 57.75 49.25 62.75
DMPO Diff-DPO 52.75 64.00 55.00 51.00 56.50
Parti Prompt DMPO SD3.5-M 63.45 71.29 63.23 52.87 67.54
Diff-DPO SD3.5-M 59.11 66.49 62.64 50.96 64.30
DMPO Diff-DPO 55.37 58.92 55.74 49.24 56.48
@@
C.3 Training Stability
To assess the stability of our optimization procedure, we examine the training loss under two different $\alpha$ values, as shown in Figure˜9. In both cases, the loss decreases smoothly and converges within a narrow range, demonstrating stable and reliable optimization dynamics throughout training.
Figure 9: Training loss curves for $\alpha=10^{-2}$ and $\alpha=10^{-3}$. In both configurations, the loss exhibits smooth decay and stable convergence, confirming the robustness of the training process.
C.4 Ablation Study
We conduct ablation studies with $\alpha$ and $\beta$ to understand the sensitivity and behavior of our alignment objective under a consistent setting: we used SD1.5 as the base model, fine-tuned with PickScore on the Pick-a-Pic V2 training dataset, and evaluated on both Pick-a-Pic V2 and HPS V2 using PickScore as the metric.
Figure 10: Ablation study on parameters $\alpha$ and $\beta$ in DMPO based on SD1.5, evaluated on the Pick-a-Pic V2 and HPS V2 test sets. (a) Effect of $\beta$: With $\alpha$ fixed at 0.01, model performance first increases and then decreases as $\beta$ grows. (b) Effect of $\alpha$: With $\beta$ fixed at 2000, performance also first increases and then decreases as $\alpha$ increases.
Figure˜10a illustrates the performance as $\beta$ increases. As $\beta$ decreases, the optimization objective degenerates into a pure reward function, leading to a drop in performance. Conversely, as $\beta$ increases, the KL-divergence penalty becomes overly restrictive, greatly limiting the model’s capacity to adapt. Figure˜10b illustrates the performance as $\alpha$ increases. Regarding the smoothing coefficient $\alpha$, we set $\alpha$ to be a small positive number since it represents the probability of the less preferred sample to avoid numerical instability. Therefore, we set $a\in\{0.1,0.01,0.001\}$ for ablation study. As shown in Appendix˜A, we observe that as $\alpha$ decreases, the second derivative of the objective $f(u)$ in Equation˜12 increases for $u>0$, resulting in a looser upper bound. This may weaken alignment precision. In particular, when $\alpha$ is very small, the reward term becomes smoothed by the coefficient $\alpha$, causing the model to occasionally favor less-preferred samples in preference pairs.